Good News and Good News

From: J. H. Lee (jhlee@sgs1.hirg.bnl.gov)
Date: Fri Aug 04 2000 - 21:07:38 EDT

  • Next message: http account: "Shift report 20000805 16:00-24:00"

     Hello All,
    
     After many "working on injection" days,  RHIC stored and cogged 
     the beam at gamma=70 for an hour this evening.  Two good news
     from the data (run# 2050)!
    
     The "vertex" distributions from Beam-Beam and ZDC seem to be
     back to normal:  time(left-right) ~ 0! After all, the shifts in
     time differences we've been seeing lately must be from actual
     changes of RHIC beam tunes, not from some extra delay cables or
     some software offsets. 
     (Please see http://pii3.brahms.bnl.gov/~operator/pictures/new_timing.ps)
    
     The completely installed Si detector seems to be working well. 
     Very nice correlations are seen between Tile and Si (ADC Sum). 
     (please see http://pii3.brahms.gov/~operator/pictures/si_vs_tile.ps).
    
     More good stable beam is expected for the weekend with hopefully
     more good news.
    
     Cheers, JH
    



    This archive was generated by hypermail 2b29 : Fri Aug 04 2000 - 21:07:12 EDT